How much energy is stored in the 180 μ𝐹 capacitor of a camera flash unit charged to

300.0 𝑉?


1
Expert's answer
2021-11-07T20:00:04-0500

"Enery, E=\frac{1}{2}CV^2=\frac{1}{2}\times 180 \times 10^{-6} \times 300^2=8.1J"